OpenStudy (anonymous):

Dumbest question I've asked in a while but I cannot remember how to multiply these (t-2t)^(1/2)*(t)^(1/2)

OpenStudy (anonymous):

because you cannot break ti up into t^1/2 and -2t^1/2 it confuses me

OpenStudy (anonymous):

there should be no way to reduce them but this solution is telling me that there is

OpenStudy (loser66):

t-2t ?

OpenStudy (anonymous):

gah flutter mistype (t-2x)*

OpenStudy (loser66):



OpenStudy (anonymous):

not t-(2xt)^(1/2) correct?

OpenStudy (loser66):

use draw box below to write it out, please.

OpenStudy (anonymous):

\[\sqrt{t-2x} \times \sqrt{t}\]

OpenStudy (loser66):